Suppose that you are looking for a student at your college who lives within five miles of you. You know that 50% of the 228 students do live within five miles of you. You randomly contact students from the college until one says he or she lives within five miles of you. Find the proability it will take

a. 6 of them until one says he or she lives within five miles of you. ____
b. 9 of them until one says he or she lives within five miles of you. ____
c. 14 of them until one says he or she lives within five miles of you. ____
Answer #1

Answer)

Probability of success here is 0.5

P = 0.5

And probability of failure (q) = 1-0.5 = 0.5

A)

We need 6th to be the probability of success

So, first five would take probability of failure

q*q*q*q*q*p

= 0.5^6

= 0.015625

B)

9

Here 9th value would be probability of success

q*q*q*q*q*q*q*q*p

= 0.001953125

C)

14

q^13*p

= 0.00006103515625